# Process Supervision-Guided Policy Optimization for Code Generation

### v1

@@ 4.3 Main Results and Analysis
Comparing Different Strategies of Using PRM in RL Training.
We explore three strategies for integrating the PRM into RL training, as described in Section 4.1: DenseReward, ValueInit, and a combined approach of DenseReward & ValueInit. The performance of RL models trained using these strategies on LiveCodeBench and InHouseBench, in comparison to SFT and RL baselines, is summarized in Table 1. For further comparison, we also include results from several publicly available models OpenAI (2023); Bai et al. (2023); Reid et al. (2024); Guo et al. (2024). Our experimental results reveal that using PRM solely as dense rewards significantly outperforms the RL baseline, consistent with findings from Wang et al. (2024a). This suggests that the granular feedback provided by PRM helps the policy explore more promising solutions by offering continuous corrections at intermediate steps. Moreover, we observe that combining PRM as both dense rewards and value function initialization results in substantial performance gains, with a $5.7\%$ improvement on LiveCodeBench and a $12.6\%$ improvement on InHouseBench, relative to the RL baseline.
Interestingly, using PRM solely for value function initialization does not provide notable benefits. We hypothesize that while value function initialization can enhance credit assignment and stabilize the learning process, it does not address the underlying issue of sparse reward signals. Without dense feedback, the policy may fail to explore the solution space effectively, resulting in limited improvement. In contrast, the combination of DenseReward and ValueInit offers a synergistic effect: dense rewards enhance exploration by providing rich intermediate feedback, while value function initialization improves stability and credit assignment. Together, these mechanisms enable the policy to converge more efficiently toward optimal solutions, which explains the significant performance improvements we observed.
PRM enhances code generation in long-horizon scenarios.
To better understand when PRM benefits code generation most, we analyze its effect based on the length of the generated responses. Intuitively, the dense nature of the reward signals provided by PRM is particularly advantageous for long-horizon tasks, where intermediate feedback can guide policy exploration more effectively. To validate this, we quantitatively compared the pass rate (Pass@1) of models trained with and without PRM across different response lengths. The results are visualized in Figure 3. Overall, the model trained with PRM demonstrates a $9\%$ improvement in Pass@1 compared to the baseline model trained without PRM. Notably, PRM consistently improves Pass@1 for responses longer than 100 tokens. However, for responses shorter than 100 tokens, PRM yields comparable or slightly inferior results. We hypothesize that in short-horizon scenarios, PRM may function similarly to a biased ORM, which limits its ability to provide meaningful improvements. Shorter responses inherently benefit less from the dense feedback signals PRM offers, as these tasks may already be well-explored by the policy without needing extensive guidance. In contrast, for more complex, long-horizon tasks, PRM offers valuable intermediate signals that effectively guide the policy to explore better solutions. These signals provide a more nuanced understanding of the correctness of individual code lines, helping the policy navigate the larger solution space with the same amount of optimization compute.
@@
Strategy LCB IHB
Full 26.9 34.6
Remove Hard 27.8 33.8
Medium Only 26.9 32.5
Revised Only 29.8 35.8
Table 2: Comparison of different PRM data selection strategies on two datasets: LCB (LiveCodeBench) and IHB (InHouseBench).
@@
Figure 4: Pass@1 on LiveCodeBench as the average number of responses per prompt for PRM data collection increases (logarithmic scale).
How much data needed to train a PRM that benefits RL training?
Given that automatic PRM data collection is computationally expensive, we examine how the performance of policies trained with PRM scales with the number of training samples. Figure 4 shows how the pass rate of models trained with varying amounts of PRM data changes along the average number of responses collected per prompt for PRM data collection, as described in Section 3.1.1. The key finding is that the performance of models trained with PRM improves consistently as the number of PRM training samples increases, highlighting the effectiveness and scalability of our approach.
@@ 5.1 LLMs for Code Generation
Recently, large language models (LLMs) have demonstrated impressive capabilities in code generation by pre-training on vast text datasets that include code (Lu et al., 2021; Christopoulou et al., 2022; Allal et al., 2023; Zheng et al., 2024; Li et al., 2023b). Additionally, models fine-tuned through supervised fine-tuning (SFT) have achieved competitive results in code generation tasks (Chen et al., 2021; Li et al., 2023a; Luo et al., 2023; Rozière et al., 2024; Guo et al., 2024). Reinforcement Learning (RL) optimizes policies by interacting with an environment and receiving rewards (Williams, 1992). Recently, RL has been incorporated into LLMs to enhance code generation using unit test feedback (Shojaee et al., 2023; Liu et al., 2023; Le et al., 2022). CodeRL (Le et al., 2022) applies unit test signals as rewards with an actor-critic method, while PPOCoder (Shojaee et al., 2023) builds on this by using the PPO algorithm. RLTF (Liu et al., 2023) improves precision by locating errors, though the reward space remains sparse. Despite progress, RL’s potential to significantly boost code generation in sparse reward environments remains underexplored.
@@ 5.2 Process Reward Models
Process reward models (PRMs) have garnered significant attention in recent LLM developments, particularly in the mathematical reasoning domain, where they provide verification for intermediate reasoning steps (Lightman et al., 2023; Wang et al., 2024a; Jiao et al., 2024; Wang et al., 2024b; Luo et al., 2024). While some approaches rely on costly and resource-intensive human-annotated process data (Lightman et al., 2023), recent research has focused on automating the collection of process supervision data Wang et al. (2024a); Jiao et al. (2024); Wang et al. (2024b); Luo et al. (2024). Building on these efforts, we similarly automate process supervision but differ in our primary objective. Rather than using PRMs solely as enhanced verifiers compared to Outcome Reward Models (ORMs), we focus on their integration into RL training for code generation. While Wang et al. (2024a) provides preliminary results on PRMs improving RL training in the mathematical domain, their findings are limited. Our work offers a more thorough and systematic investigation of how PRMs can be leveraged in RL for code generation tasks.
@@ 6 Conclusions and Limitations
In this work, we addressed the challenge of sparse reward signals in reinforcement learning (RL) for code generation by introducing a Process Reward Model (PRM) that provides dense, line-level feedback. This approach mirrors human-like code refinement and improves learning efficiency. Our experiments demonstrate that integrating PRMs significantly enhances the pass rates of code generation models on both the LiveCodeBench dataset and proprietary benchmarks. This method has the potential to improve long-horizon code generation scenarios, advancing the state-of-the-art in LLM-based code generation.
@@
Despite the promising results, our approach has several limitations that warrant further investigation. First, the effectiveness of the PRM relies heavily on the quality of the collected data. While we automated data collection using binary search and unit tests, this method may not capture all nuances of code correctness and may introduce noise, especially in more complex or ambiguous programming tasks. Second, the computational cost of collecting PRM training data is still substantial, even though we employed binary search to mitigate it. Third, our automated PRM data collection method requires external verification (unit tests in our case), which is not applicable to many other domains such as creative writing or open-ended generation problems. This could limit the applicability of our approach in those areas.
Appendix A Appendix
@@
Figure 5: RL training curve of our method (DenseReward&ValueInit) compared to other settings. Using PRM as both DenseReward and ValueInit yields the best result.

### v2

@@ 4.3 Main Results and Analysis
Comparing PRM Integration Strategies in RL Training.
We evaluate three strategies for incorporating PRM into RL training, as outlined in Section 4.1: DenseReward, ValueInit, and a combined approach (DenseReward & ValueInit). Table 1 presents the performance of RL models trained using these strategies on HumanEval, MBPP, and LiveCodeBench, alongside SFT and RL baselines. For reference, we also include results from publicly available models such as GPT-4o-mini (OpenAI, 2023), Qwen2-72B (Bai et al., 2023), and DeepseekCoder-33B (Guo et al., 2024).
Our results indicate that using PRM as a dense reward signal significantly improves performance over the RL baseline (see the first and third settings for Qwen2.5-7B-RL and Doubao-Lite-RL in Table 1), aligning with findings from (Wang et al., 2024a). The granular feedback provided by PRM facilitates policy exploration by offering continuous corrections at intermediate steps. Additionally, using PRM solely for value function initialization also yields consistent improvements. In Qwen2.5-7B-RL, this setting outperforms the RL baseline across all benchmarks, while in Doubao-Lite-RL, similar gains are observed on HumanEval and MBPP.
Combining PRM for both dense rewards and value initialization yields significant performance improvements. In Qwen2.5-7B-RL, Pass@1 increases from $62.4\%$ to $65.4\%$ on MBPP and from $27.5\%$ to $30.1\%$ on LiveCodeBench. Similarly, in Doubao-Lite-RL, Pass@1 improves from $65.1\%$ to $70.9\%$ on HumanEval, $61.9\%$ to $63.8\%$ on MBPP, and $28.2\%$ to $29.8\%$ on LiveCodeBench. This improvement stems from the complementary roles of PRM: dense rewards facilitate exploration by providing rich intermediate feedback, while value initialization stabilizes training and enhances credit assignment. Together, these mechanisms accelerate convergence toward optimal solutions, driving the observed performance gains.
PRM Encourages Exploration and Improves Learning Efficiency.
We evaluated the Best-of-K performance for all four training configurations in Doubao-Lite-RL experiments on the training set. Specifically, we assessed all RL models using a decoding configuration with a temperature of 1.0, nucleus sampling (top-$p=0.95$), and top-k sampling ($k=128$). For each $K$, we recorded the percentage of problems solved within $K$ generated responses, referred to as the Pass Rate. Figure 3 presents the results for $K$ ranging from 1 to 30. Both DenseReward and ValueInit individually enhance Best-of-K performance compared to the baseline. When combined, they yield the highest improvement, with a nearly 4% increase in Pass Rate at $K=30$ over the baseline, highlighting the synergy between dense rewards and value initialization.
Figure 3: Best-of-K performance curves for all RL training settings, showing the percentage of problems solved within $K$ generated responses.
PRM Enhances Code Generation in Long-Horizon Scenarios.
To understand when PRM provides the greatest benefit, we analyze its impact based on response length. Intuitively, the dense nature of the reward signals provided by PRM is particularly advantageous for long-horizon problems, where intermediate feedback can guide policy exploration more effectively.
To validate this, we compared Pass@1 performance of models trained with and without PRM across different response lengths, as shown in Figure 4. Overall, PRM-trained models achieve a 9% improvement in Pass@1 over the baseline. Notably, PRM consistently enhances performance for responses exceeding 100 tokens, whereas for shorter responses, its effect is neutral or slightly negative. We hypothesize that in short-horizon problems, PRM behaves similarly to a biased ORM, offering limited improvements since these problems are already well-explored by the policy. In contrast, for complex, long-horizon problems, PRM provides valuable intermediate signals that help the policy navigate the solution space more effectively, achieving better results with the same optimization compute.
@@
Table 2: Comparison of different PRM data selection strategies on LiveCodeBench (LCB) with Doubao-Lite-RL models.
Full Remove Hard Medium Only Revised Only
LCB 26.9 27.8 26.9 29.8
How much data needed to train a PRM that benefits RL training?
Given that automatic PRM data collection is computationally expensive, we examine how the performance of policies trained with PRM scales with the number of training samples. Figure 5 shows how the pass rate of Doubao-Lite-RL models trained with varying amounts of PRM data changes along the average number of responses collected per prompt for PRM data collection, as described in Section 3.1.1. The key finding is that the performance of models trained with PRM improves consistently as the number of PRM training samples increases, highlighting the effectiveness and scalability of our approach.
@@
Figure 5: Pass@1 on LiveCodeBench as the average number of responses per prompt for PRM data collection increases (logarithmic scale). A value of $<2^{0}$ indicates that we subsampled prompts from the full dataset, resulting in a smaller prompt set.
@@ 5.1 LLMs for Code Generation
Recently, large language models (LLMs) have demonstrated impressive capabilities in code generation by pre-training on vast text datasets that include code (Lu et al., 2021; Christopoulou et al., 2022; Allal et al., 2023; Zheng et al., 2024; Li et al., 2023b). Additionally, models fine-tuned through supervised fine-tuning (SFT) have achieved competitive results in code generation tasks (Chen et al., 2021a; Li et al., 2023a; Luo et al., 2023; Rozière et al., 2024; Guo et al., 2024). Reinforcement Learning (RL) optimizes policies by interacting with an environment and receiving rewards (Williams, 1992). Recently, RL has been incorporated into LLMs to enhance code generation using unit test feedback (Shojaee et al., 2023; Liu et al., 2023; Le et al., 2022). CodeRL (Le et al., 2022) applies unit test signals as rewards with an actor-critic method, while PPOCoder (Shojaee et al., 2023) builds on this by using the PPO algorithm. RLTF (Liu et al., 2023) improves precision by locating errors, though the reward space remains sparse. Despite progress, RL’s potential to significantly boost code generation in sparse reward environments remains underexplored.
@@ 5.2 Process Reward Models
Process reward models (PRMs) have garnered significant attention in recent LLM developments, particularly in the mathematical reasoning domain, where they provide verification for intermediate reasoning steps (Lightman et al., 2023; Wang et al., 2024a; Jiao et al., 2024; Wang et al., 2024b; Luo et al., 2024). While some approaches rely on costly and resource-intensive human-annotated process data (Lightman et al., 2023), recent research has focused on automating the collection of process supervision data (Wang et al., 2024a; Jiao et al., 2024; Wang et al., 2024b; Luo et al., 2024). Building on these efforts, we similarly automate process supervision but differ in our primary objective. Rather than using PRMs solely as enhanced verifiers compared to Outcome Reward Models (ORMs), we focus on their integration into RL training for code generation. While (Wang et al., 2024a) provides preliminary results on PRMs improving RL training in the mathematical domain, their findings are limited. Our work offers a more thorough and systematic investigation of how PRMs can be leveraged in RL for code generation tasks.
@@ 6 Conclusions and Limitations
In this work, we tackled the challenge of sparse reward signals in RL for code generation by introducing a PRM that provides dense, line-level feedback. This approach, inspired by human-like code refinement, enhances learning efficiency and encourages better exploration. Our experiments show that integrating PRMs significantly improves the pass rates of code generation models across HumanEval, MBPP, and LiveCodeBench. Notably, PRM not only facilitates more effective RL training but also improves performance in long-horizon code generation scenarios.
@@
Despite these promising results, our approach has several limitations that warrant further exploration. First, PRM effectiveness depends on the quality of the collected data. While we automate data collection using binary search and unit tests, this method may overlook nuances of code correctness and introduce noise, particularly in complex or ambiguous programming tasks. Second, despite using binary search to reduce overhead, PRM training remains computationally expensive. Third, our method relies on external verification (e.g., unit tests), which limits its applicability to domains lacking well-defined correctness criteria, such as creative writing or open-ended generation tasks. Addressing these challenges presents exciting future research directions including improving PRM data collection strategies and exploring alternative evaluation methods to extend PRM applicability beyond structured domains like code generation.
Appendix A A Typical Example of the Learned Line-wise Rewards
In Figure 6, we present a typical example of the line-wise rewards identified by binary search and predicted by a learned PRM to give readers a clearer understanding of our method. In this example, we first sampled a problem from the training set and used our in-house model to generate a response for it. For this generated response (which is not included in the PRM training data), we show the line-wise rewards derived from two sources:
  1. 1.
    Line-wise Rewards Identified by Binary Search: We directly applied the model to perform Algorithm 1, labeling the reward for each line.
  2. 2.
    Line-wise Rewards Predicted by a Learned PRM: We used the learned PRM to predict the rewards for each line.
Line-wise Rewards Identified by Binary Search Line-wise Rewards Predicted by a Learned PRM
Refer to caption Refer to caption
Figure 6: Visualization of the learned line-wise rewards. The top gray block displays the problem description, while the bottom section shows a model-generated response with line-wise rewards from different sources. The bottom-left block presents the line-wise rewards identified by binary search, and the bottom-right block presents the line-wise rewards predicted by a learned PRM. The actual reward value is shown at the beginning of each line, and each line is color-coded based on the reward value: lines with rewards closer to -1 are shaded red, while those closer to +1 are shaded green.
Appendix B RL Training Curves
In Figure 7, we present the smoothed RL training curves for all four settings (with and without DenseReward, and with and without ValueInit) in Doubao-Lite-RL experiments, using a moving average to reduce noise and enhance readability. These curves correspond to all four RL settings reported in Table 1. The smoothed trends clearly show that when PRM is used as DenseReward, the model solves more problems compared to the baseline, demonstrating PRM’s role in enabling more efficient exploration during RL training. Furthermore, when PRM is applied as both DenseReward and ValueInit, our method achieves the best performance.
@@ Appendix B RL Training Curves
Figure 7: RL training curve of all experiment settings in Doubao-Lite-RL experiments. Using PRM as both DenseReward and ValueInit (DenseReward&ValueInit) yields the best result.
Appendix C PRM Training Data Statistics
We present detailed statistics on the PRM datasets used in our experiments to evaluate the optimal PRM data selection strategy, as discussed in Section 4.3. The experiments were conducted using the Doubao-Lite series models. Table 3 summarizes the following key metrics: the number of prompt-response pairs (#Samples); the total number of tokens across all responses (#Tokens); the average number of lines in all responses (Avg. #Lines); and the distribution of PRM labels (-1/0/+1).
In Figure 8, we present the distribution of error positions of all Revised responses (responses that initially fail but have a correct prefix identified) as determined by the Binary Search procedure (Algorithm 1). The absolute error position (i.e., the position of the first token rejected by Binary Search) is normalized as follows: for a response $\mathbf{y}=(y_{1},y_{2},\ldots,y_{L})$ with $L$ tokens, if the Binary Search accepted the prefix $(y_{1},y_{2},\ldots,y_{p})$ consisting of $p$ tokens, the Relative Error Position is calculated as $\frac{p}{L}$.
Table 3: Statistics of PRM training data collected using different data selection strategies.
Strategy #Samples #Tokens Avg. #Lines PRM Labels
-1 0 +1
Full 838K 179M 16.82 44.25% 17.87% 37.88%
Remove Hard 630K 119M 15.06 24.03% 19.18% 56.79%
Medium Only 485K 104M 16.57 27.66% 19.41% 52.93%
Revised Only 352K 76M 16.71 13.20% 19.42% 67.38%
Figure 8: Distribution of Relative Error Positions Identified by Binary Search.
Appendix D Qwen2.5-7B Experiment Details
Base Model.
We adopt Qwen2.5-7B as our base model (QwenTeam, 2024), a recently released causal language model available at https://huggingface.co/Qwen/Qwen2.5-7B. Qwen2.5 belongs to the Qwen series of large language models (Yang et al., 2024), known for their advanced capabilities across a wide range of domains. The Qwen2.5-7B model has 7.61 billion total parameters (6.53 billion excluding embeddings) and utilizes the Transformer architecture as its core. It incorporates state-of-the-art enhancements, including Rotary Positional Embedding (RoPE), SwiGLU activation, RMSNorm, and Attention QKV bias. The model consists of 28 layers and employs 28 attention heads for queries (Q) and 4 for keys and values (KV), making it highly efficient for tasks requiring robust attention mechanisms.
SFT Settings.
We fine-tuned the Qwen2.5-7B model on the SFT dataset described in Section 4.1. The model was trained for two epochs, starting with a learning rate of $1\times 10^{-7}$, which linearly increased to $2\times 10^{-5}$ during the first 2% of the total training steps. After reaching the peak learning rate, a cosine learning rate decay schedule was applied, gradually reducing the learning rate to $2\times 10^{-6}$ for the remainder of the training. Additionally, a constant weight decay of 0.01 was used throughout the SFT training process to regularize the model and improve generalization. The model fine-tuned through this process is referred to as Qwen2.5-7B-SFT.
RL Baseline.
We adopted the same RL baseline training method and used the same RLHF dataset described in Section 4.1 to further train the Qwen2.5-7B-SFT model. For PPO training, we configured the following hyperparameters: a batch size of 4096, a linear warmup over the first 5 steps, followed by a constant learning rate of $2\times 10^{-6}$ for both the actor and critic, and a KL penalty of 0.01. The training utilized the AdamW optimizer and spanned approximately 300 steps, during which we empirically observed performance convergence.
PRM Training.
We selected four checkpoints at 50, 100, 150, and 200 steps during the training process of the RL baseline model. For each checkpoint, we sampled $n=5$ responses for every coding prompt in the training dataset $\mathcal{D}_{\text{train}}$. Each sampled response was labeled using the binary search procedure described in Algorithm 1, with $K=20$ completions generated for each partial code prefix. The data collected from all checkpoints was then aggregated to form a PRM training set, employing the Revised Only strategy described in Section 4.3. This resulted in 165K samples and 28M tokens. On average, each response contained 16.07 lines. The PRM label distribution was 25.88% for -1, 15.90% for 0, and 58.22% for +1. The PRM was initialized using the value model from the RL baseline and fine-tuned on this PRM dataset using the objective function defined in (2).
Integrating PRM into RL.
[…]
{198, 510, 982, 340, 271, 2398, 921, 741, 3932, 1171, 692, 1305, 4167, 2546, 1447, 10343, 1138, 19324,
341, 5563, 9957, 382, 3407, 3646, 624, 48443, 280, 456, 2533, 3989, 1248, 5613, 8389, 8997, 698,
24135, 317, 7368, 2440, 10907, 22165, 4432, 5929, 7129, 345, 11043, 532, 4660, 21686, 14288}.
During RL training, we only applied partial rewards from PRM to these tokens.